In a certain small town, Comcast has 100000 subscribers. According to Census data, 30 percent of the subscribers are single parent households. A simple random sample of 400 households were selected for a survey. The percentage of single parent households in the sample will be around ______, give or take______ or so. Estimate the chance that between 25.4 percent and 34.6 percent of the households in the sample are single parent households.___________
Answer #1
std error of proportion=σp=√(p*(1-p)/n)= 0.0229

The percentage of single parent households in the sample will be around 30 % give or take 2.29 %

probability = P(0.254<X<0.346) = P(-2.01<Z<2.01)= 0.9778-0.0222= 0.9556